OpenStudy (anonymous):

(-14) + 81 = 81 + (-14) associative property of addition distributive property multiplicative inverse additive identity commutative property of addition Honestly don't know this stuff.

OpenStudy (bahrom7893):

Which basically means that a+b=b+a

OpenStudy (bahrom7893):

You should look up all the other properties, so that you know next time :)

OpenStudy (anonymous):

thanks.(:

OpenStudy (bahrom7893):

associative is (a+b)+c = a+(b+c)

OpenStudy (anonymous):

101 + (29 + 417) = (101 + 29) + 417 so this is associative ?

OpenStudy (bahrom7893):

yup

OpenStudy (anonymous):

cool